OpenStudy (anonymous):

Showing all of your work, solve for x: 7x - 3 = 2(x + 6)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

7x-3=2x+6 7x=2x+3 5x=3 x=1.6666666666

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation. 7x−3=2(x+6) Simplify: (Show steps) 7x−3=2x+12 Step 2: Subtract 2x from both sides. 7x−3−2x=2x+12−2x 5x−3=12 Step 3: Add 3 to both sides. 5x−3+3=12+3 5x=15 Step 4: Divide both sides by 5. 5x/5 =15/5 =3 x=3 Answer: x=3

OpenStudy (anonymous):

oops i didnt times my 6 by 2 sorry the correct awnser is x=3

OpenStudy (anonymous):

...